A wire fixed at the upper end stretches by length l by applying a force F. The work done in stretching is
(a) 2F l
(b) f l
(c) F/2 l
(d) F l/2

Grade:12th pass

1 Answers

Aditi Chauhan
9 years ago

Sol. 2. (d). Work done by constant force in displacing the object by a distatance l

= 1/2 * stress * strain * volume

= 1/2 / F/A * l/L * A * L = Fl/2